Show all the solutions. Round off answers to two decimal places. Be consistent with the units.

1, Given with the accepted value of the resistivity for copper, solve for the percentage error.  The average resistivity is the experimental value.

         ρ (copper) = 1.72 x 10-6 Ω-cm






1
Expert's answer
2021-04-21T13:14:16-0400

We have given the expected value or the average value of resistivity "= 1.72\times 10^{-6} \Omega m"

We know,

"Percentage\hspace{1mm}error = \dfrac{Actual\hspace{1mm} value-Expected \hspace{1mm}value}{Expected \hspace{1mm}value}100" %


Hence, we can assume that the actual value will be "= 1.74 \times 10^{-6}"



Hence Percentage error = "\dfrac{1.74-1.72}{1.72}\times 100" %


"= 0.28%" %